# Double-Signed Fragmented DNSSEC for Countering Quantum Threat

### v1

@@ 5. Evaluation Setup
For experimental purposes, our ‘client’ is a Ubuntu OS (22.04) that we simply use for querying our resolver and for analyzing the DNS response message. All of the other components - i.e., ‘Resolver’, ‘Root Server’, and ‘Authoritative’ servers run instances of Internet System Consortium’s BIND9 as the DNS software (see details of BIND9 below). We create a simple ‘Docker Network’ (using Docker Version 25.0.3) to run all of the aforementioned components in a separate Docker container on an Amazon EC2 t2.xlarge instance which provides 4 cores of 2.3GHz Inter Xeon CPU and 16GB of RAM instance. Since by default, BIND9 does not support NIST’s quantum-safe digital signature algorithms, we adopt the prototype implementation of OQS-BIND from (G, 2021) that adds the support of FALCON512, DILITHIUM2-AES, and SPHINCS+-SHA256-128S (with parameters set for level 1 - i.e.. 128-bit-security) using OpenSSL 1.1.1 the Open Quantum Safe (OQS) liboqs library 0.7.2 (Stebila and Mosca, 2017).
@@
  • Step 1 Our client (a simple docker container with Ubuntu OS) wants to resolve ‘test.socratescrc’ (note that, for the experimental purpose we have assigned an ‘A’ RR to this subdomain on our authoritative server). It queries the ‘Resolver’ for address resolution.
@@
To incorporate the fragmentation strategy depicted in Figure 10 in this evaluation setup, we utilized the QBF daemon since it adopts the standard RR format unlike ARRF that relies on custom-built RRs (Rawat and Jhanwar, 2024). We included all the steps needed for accommodating double signatures in the daemon. This daemon then intercepts all the incoming and outgoing DNS packets on the resolver, root server, and name server and performs the relevant steps as needed (see details in Section 4 and Fig 10). Therefore, practically, each component - i.e., resolver, root server, and authoritative server in our evaluation setup has an OQS-BIND software and daemon that incorporate our fragmentation strategy for double signatures (see overview in Figure 12). On the name servers, the daemon intercepts the incoming query, updates the maximum advertised UDP size to 64KB (i.e., max possible), and forwards it to the OQS-BIND software. The OQS-BIND creates a DNSSEC response to this query with double signatures, which is again intercepted by our daemon that executes all the steps depicted in Fig. 10. Upon receiving the response from the name server, the resolver’s daemon intercepts it and goes through all the steps needed for accommodating and re-assembling double signatures (see Section 4 and Fig. 10). Once the resolver’s daemon receives all the fragments, it reassembles them and finally sends the reassembled message to the OQS-BIND, which initiates the verification process for both signatures.
@@
Berkely Internet Name Domain (BIND)
BIND offers a complete suite of the DNS protocol, and the BIND9 software distribution contains a name server (named) and a set of associated tools such as for key and signature generation that are needed for our analysis. It is one of the most widely used software on DNS servers on the Internet (Knudsen, 2021). Since BIND9 can be configured (using named.conf) as a resolver or name server, it allowed us to run ‘Resolver’, ‘Root Server’, and ‘Authoritative Server’ in separate containers using the same BIND9 docker image. Although BIND9 allows flexible configuration features and supports multiple functions such as a single DNS server acting as both an authoritative name server and a resolver, we in our testbed have assigned one dedicated container to the resolver and authoritative name server as typically preferred by large operators (Manual, 2023). Therefore, our testbed represents a realistic small-scale DNS infrastructure for assessing the impact of quantum-safe algorithms.
@@
Our evaluation setup as described above is capable of performing the address resolution steps of conventional DNSSEC with both pre-quantum and post-quantum digital signatures. However, since we want the support of double signatures added to our testbed, we conducted experiments to ascertain whether BIND9 by default supports two signatures. Our observations are as follows:
Signature Generation on Name Servers: Our analysis revealed that we can create two key pairs and signatures - i.e., both pre-quantum and post-quantum on name servers (i.e., root and authoritative servers) by using the BIND9 utilities (see dnssec-keygen and dnssec-signzone tools in BIND9 manuals). Both signatures can be added directly to the corresponding signed zone files generated in BIND9. Figure. 13 depicts a portion of the signed zone file generated on the authoritative server. It is evident that the signed zone file contains two signatures for a representative record -i.e., test0.socratescrc..
@@
Double Signature Verification on Resolver: Our analysis of the verification process revealed that although we can send both signatures to the resolver and re-assemble them correctly in our test setup, the BIND9 signature validation by default checks only one signature before marking any response from the name server(s) as secure. Therefore, we modified the BIND9 source code (especially the code corresponding to the validation of the received responses) so that both the re-assembled signatures must be valid before accepting any response. Fig. 14 depicts a portion of the log file of our modified resolver. It is evident that the modified resolver in our setup verifies both signatures, unlike the default BIND9 resolver.
@@
For analyzing the impact of double signatures in the DNSSEC address resolution process, we kept 10 A resource records with unique labels on our authoritative server (e.g., test0.socratescrc – test10.socratescrc). Then, using the BIND9 key and signature generation tools, we generated two signatures with various possible combinations of pre-quantum (ECDSA256, RSASHA256) and post-quantum (FALCON512, DILITHIUM2, SPHINCS+-SHA256-128S) digital signature algorithms. Once both signatures were generated, we moved the DS RR of the authoritative server to the root server and the DS RR of the root server to the resolver as a trust anchor so that we could establish the chain of trust identical to the conventional DNSSEC. Once all of this process is complete, using our client, we query the resolver iteratively for 10 resource records that we kept on our authoritative serve in our test setup. For each query, we estimate the query resolution time under simulated network conditions of bandwidth of 50Mbps and delay of 10ms (note that, we use the Linux tc utility for simulating these bandwidth and latency values). Finally, we compare the average resolution time of double signatures with post-quantum signatures to find any issues in migrating from only post-quantum signatures to double signatures for the interim period until CRQCs are fully realized as discussed in Section 1. A similar experimental approach is also used in (Rawat and Jhanwar, 2024) but with only one signature for the post-quantum era. Additionally, we also investigate how many additional fragments our daemon needs to compute and re-assemble with double signatures in comparison with post-quantum signatures alone.
@@ 6. Experimental Results
DILITHIUM+ECDSA 23 15
DILITHIUM+RSA 23 15
@@
In this section, we present the results of our experiments described in the previous section. Our thorough analysis suggests that our approach successfully incorporates all possible combinations of double signatures (pre-quantum + post-quantum) in DNSSEC, fragments them appropriately on name servers, and re-assembles the responses correctly on the resolver side. Additionally, we also observed that all the double signatures were correctly verified on the resolver. Table 1 shows the number of fragments when various combinations of pre- and post-quantum signatures are employed in our DNSSEC setup. Analysis suggests that our approach only adds one additional fragment in response to QTYPE ‘A’ and ‘DNSKEY’ queries when FALCON and DILITHIUM2 are combined with any of the pre-quantum signatures - i.e., ECDSA256 and RSASHA256 in comparison with post-quantum signatures alone. Furthermore, we observed that combining SPHINCS+ with any of the pre-quantum signatures does not add additional fragments. This is due to the reason that after sending the initial fragment, the daemon on name servers removes all RRs except RRSIG, DNSKEY, and OPT, thereby accommodating double-signatures created with SPHINCS+ in same number of fragments as with SPHINC+ alone. Mean resolution times with various combinations of double-signatures are shown in Tables 2, 3, 4. The results show that the impact of double signatures is insignificant on the resolution times of queries when compared with post-quantum signatures alone. In any double-signature combination that we have tested, resolution time only increases by less than 9% as compared with the post-quantum signatures alone. For instance, the resolution time for a combination of FALCON + ECDSA is 205.9ms as compared with 190.1ms for FALCON alone (i.e., an increase of approximately 8.3%). The lowest resolution time for double signatures was observed for any combination of FALCON with either RSA or ECDSA-based pre-quantum digital signatures. This is because FALCON has the shortest signature size compared with other post-quantum digital signatures. Similarly, RSA combined with any of three post-quantum candidates results in a lower resolution time as compared with ECDSA (e.g., DILITHIUM + RSA results in lower resolution time in comparison with DILITHIUM + ECDSA, and likewise for other two post-quantum digital signatures). This is because RSA has faster signature verification than ECDSA (Brady, 2020).
@@
The aforementioned observations suggest that double signatures can reliably be incorporated into DNNSEC infrastructure and have a minimal impact on performance as compared with post-quantum digital signatures alone. Therefore, their incorporation is ideally suited for DNSSEC during the interim period from now until CRQCs are fully realized. Particularly, the double signatures created with either FALCON or DILITHIUM with RSA or ECDSA have a resolution time that is 14-22% less than that of SPHINCS+ alone. Since both FALCON and DILITHIUM are based upon lattices, SPHINCS+ was only added as a post-quantum candidate by NIST to avoid the over-reliance on the lattices (see (NIST, 2022)). However, as observed in our experiments, SPHINCS+ has a significantly large signature size that results in much higher resolution time in DNSSEC. Therefore, the double-signature created with either FALCON or DILITHIUM is the optimal solution for the specified interim period while avoiding the envisaged problems of SPHINCS+ (see results described above and depicted in Tables 2 - 4). These double-signatures will essentially offer additional security to DNSSEC during the interim period if any attacks on lattice-based signatures are revealed in the near future that could be perpetrated through today’s classical computer.
@@ 7. Related Work
Our literature review suggests that, historically there have been a few proposals for dealing with large message size issues. For example, authors in (M. Sivaraman, 2015; Song and Wang., 2019) proposed potential ways for conducting fragmentation at the application layer. However, both of them can potentially lead to DoS amplification attacks since they send many packets in response to a single query. Additionally, there are a few recent research works that focus on transitioning toward quantum-safe DNSSEC and related issues. However, unlike this work, none of these prior research works consider double-signatures for DNSSEC. For example, authors in (Müller et al., 2020; Beernink, 2022; Goertzen and Stebila, 2022) investigated NIST Round 3 and 4 PQC candidates, respectively, for their utilization in DNSSEC. Müller et al. (Müller et al., 2020) specifically analyzed NIST Round 3 candidates and established the various requirements that different candidates either meet or do not for their incorporation in DNSSEC. Similarly, Beernink et al. (Beernink, 2022) presented a way of delivering larger keys out-of-band from DNSSEC. Goertzen et al. (Goertzen and Stebila, 2022) focused on the fragmentation issue in DNSSEC for the post-quantum era. However as discussed before, it adopts a custom-designed non-standard pseudo-RR for performing the fragmentation at the application layer. The usage of non-standard RR makes its practicability questionable. Similarly, an improved version of the fragmentation approach for quantum-safe DNSSEC is also presented in (Rawat and Jhanwar, 2024), which leverages only standard RRs and thus seems more practical. However, unlike these existing works, we investigate the possibility of using double signatures for DNSSEC, and how the fragmentation issue can be dealt with in the simultaneous presence of pre and post-quantum signatures in DNSSEC. Through our implemented testbed and empirical analysis, we show the efficacy of double signatures in DNSSEC for the interim period until CRQCs are fully realized.
@@ 8. Conclusion & Future Work
In this work, we investigated the possibility of using double signatures in DNSSEC. For this purpose, we first investigated the maximum message size issue resulting from the incorporation of double signatures in-depth and then devised an approach that can successfully incorporate double signatures on name servers and deliver them to the resolver over UDP by conducting the fragmentation of response messages at the application layer. We then built a functional Docker-based testbed by leveraging OQS-BIND9 and a daemon that handles double signatures and the fragmentation strategy. Our investigation suggests double signatures can be generated and added to signed zone files on the authoritative name servers. However, the source code of BIND9 needs a moderate level of changes to facilitate the verification of two signatures for the same resource record. With these required changes made to the BIND9 resolver and clubbing our daemon with it, we comprehensively analyzed the resolution time for various combinations of pre and post-quantum digital signatures. Our analysis revealed that double signatures do not have any significant overhead as compared with post-quantum digital signatures alone. Particularly, our investigation suggests that lattice-based signatures - i.e., FALCON and DILITHIUM combined with any of the pre-quantum digital signatures results in a shorter resolution time than SPHINCS+ alone which was shortlisted (despite having a very large public key and signature size and a slower verification) by NIST to avoid over-reliance on the lattice-based signatures for the post-quantum era. Our analysis suggests that double signatures generated with FALCON and DILITHUM are well-suited for DNSSEC and offer performance benefits compared with SPHINCS+ alone. The double signatures approach will also offer additional security in case any attacks on lattice-based PQ candidates are revealed in the near future. While the analysis conducted in this work establishes the overall efficacy of double signatures for DNSSEC, we believe that resolution time can further be reduced by incorporating our fragmentation strategy directly into BIND9 rather than in the form of a daemon. We plan to make these advancements and relevant analyses in the future.

### v2

@@ 5. Evaluation Setup
For experimental purposes, our ‘client’ is a Ubuntu OS (22.04) that we simply use for querying our resolver and for analyzing the DNS response message. All of the other components - i.e., ‘Resolver’, ‘Root Server’, and ‘Authoritative’ servers run instances of Internet System Consortium’s BIND9 as the DNS software (see details of BIND9 below). We create a simple ‘Docker Network’ (using Docker Version 25.0.3) to run all of the components as mentioned earlier in a separate Docker container on an Amazon EC2 t2.xlarge instance, which provides 4 cores of 2.3GHz Intel Xeon CPU and 16GB of RAM. Since by default, BIND9 does not support NIST’s quantum-safe digital signature algorithms, we adopt the prototype implementation of OQS-BIND from (G, 2021) that adds the support of FALCON512, DILITHIUM2-AES, and SPHINCS+-SHA256-128S (with parameters set for level 1 - i.e, 128-bit-security) using OpenSSL 1.1.1, the Open Quantum Safe (OQS) liboqs library 0.7.2 (Stebila and Mosca, 2017).
@@
  • Step 1 Our client (a simple Docker container with Ubuntu OS) wants to resolve ‘test.socratescrc’ (note that, for experimental purpose we have assigned an ‘A’ RR to this subdomain on our authoritative server). It queries the ‘Resolver’ for address resolution.
@@
To incorporate the fragmentation strategy depicted in Figure 10 in this evaluation setup, we utilized the QBF daemon since it adopts the standard RR format, unlike ARRF that relies on custom-built RRs (Rawat and Jhanwar, 2024). We included all the steps needed for accommodating double signatures in the daemon. This daemon then intercepts all the incoming and outgoing DNS packets on the resolver, root server, and name server and performs the relevant steps as needed (see details in Section 4 and Fig 10). Therefore, practically, each component - i.e., resolver, root server, and authoritative server in our evaluation setup has an OQS-BIND software and daemon that incorporates our fragmentation strategy for double signatures (see overview in Figure 12). On the name servers, the daemon intercepts the incoming query, updates the maximum advertised UDP size to 64KB (i.e., max possible), and forwards it to the OQS-BIND software. The OQS-BIND creates a DNSSEC response to this query with double signatures, which is again intercepted by our daemon that executes all the steps depicted in Fig. 10. Upon receiving the response from the name server, the resolver’s daemon intercepts it and goes through all the steps needed for accommodating and re-assembling double signatures (see Section 4 and Fig. 10). Once the resolver’s daemon receives all the fragments, it reassembles them and finally sends the reassembled message to the OQS-BIND, which initiates the verification process for both signatures.
@@
Berkeley Internet Name Domain (BIND)
BIND offers a complete suite of the DNS protocol, and the BIND9 software distribution contains a name server (named) and a set of associated tools such as for key and signature generation, that are needed for our analysis. It is one of the most widely used software on DNS servers on the Internet (Knudsen, 2021). Since BIND9 can be configured (using named.conf) as a resolver or name server, it allowed us to run ‘Resolver’, ‘Root Server’, and ‘Authoritative Server’ in separate containers using the same BIND9 docker image. Although BIND9 allows flexible configuration features and supports multiple functions such as a single DNS server acting as both an authoritative name server and a resolver, we in our testbed have assigned one dedicated container to the resolver and authoritative name server as typically preferred by large operators (Manual, 2023). Therefore, our testbed represents a realistic small-scale DNS infrastructure for assessing the impact of quantum-safe algorithms.
@@
Our evaluation setup, as described above, is capable of performing the address resolution steps of conventional DNSSEC with both pre-quantum and post-quantum digital signatures. However, since we want the support of double signatures added to our testbed, we conducted experiments to ascertain whether BIND9 by default supports two signatures. Our observations are as follows:
Signature Generation on Name Servers: Our analysis revealed that we can create two key pairs and signatures - i.e., both pre-quantum and post-quantum on name servers (i.e., root and authoritative servers) by using the BIND9 utilities (see dnssec-keygen and dnssec-signzone tools in the BIND9 manuals). Both signatures can be added directly to the corresponding signed zone files generated in BIND9. Figure. 13 depicts a portion of the signed zone file generated on the authoritative server. It is evident that the signed zone file contains two signatures for a representative record -i.e., test0.socratescrc..
@@
Double Signature Verification on Resolver: Our analysis of the verification process revealed that although we can send both signatures to the resolver and reassemble them correctly in our test setup, the BIND9 signature validation by default checks only one signature before marking any response from the name server(s) as secure. Therefore, we modified the BIND9 source code (especially the code corresponding to the validation of the received responses) so that both the re-assembled signatures must be valid before accepting any response. Fig. 14 depicts a portion of the log file of our modified resolver. It is evident that the modified resolver in our setup verifies both signatures, unlike the default BIND9 resolver.
@@
For analyzing the impact of double signatures in the DNSSEC address resolution process, we kept 10 A resource records with unique labels on our authoritative server (e.g., test0.socratescrc – test10.socratescrc). Then, using the BIND9 key and signature generation tools, we generated two signatures with various possible combinations of pre-quantum (ECDSA256, RSASHA256) and post-quantum (FALCON512, DILITHIUM2, SPHINCS+-SHA256-128S) digital signature algorithms. Once both signatures were generated, we moved the DS RR of the authoritative server to the root server and the DS RR of the root server to the resolver as a trust anchor so that we could establish the chain of trust identical to the conventional DNSSEC. Once all of this process is complete, using our client, we query the resolver iteratively for 10 resource records that we kept on our authoritative server in our test setup. For each query, we estimate the query resolution time under simulated network conditions of bandwidth of 50Mbps and delay of 10ms (note that we use the Linux tc utility for simulating these bandwidth and latency values). Finally, we compare the average resolution time of double signatures with post-quantum signatures to find any issues in migrating from only post-quantum signatures to double signatures for the interim period until CRQCs are fully realized as discussed in Section 1. A similar experimental approach is also used in (Rawat and Jhanwar, 2024), but with only one signature for the post-quantum era. Additionally, we also investigate how many additional fragments our daemon needs to compute and reassemble with double signatures in comparison with post-quantum signatures alone.
@@ 6. Experimental Results
SPHINCS+ECDSA 23 15
SPHINCS+RSA 23 15
@@
In this section, we present the results of our experiments described in the previous section. Our thorough analysis suggests that our approach successfully incorporates all possible combinations of double signatures (pre-quantum + post-quantum) in DNSSEC, fragments them appropriately on name servers, and reassembles the responses correctly on the resolver side. Additionally, we also observed that all the double signatures were correctly verified on the resolver. Table 1 shows the number of fragments when various combinations of pre- and post-quantum signatures are employed in our DNSSEC setup. Analysis suggests that our approach only adds one additional fragment in response to QTYPE ‘A’ and ‘DNSKEY’ queries when FALCON and DILITHIUM2 are combined with any of the pre-quantum signatures - i.e., ECDSA256 and RSASHA256 in comparison with post-quantum signatures alone. Furthermore, we observed that combining SPHINCS+ with any of the pre-quantum signatures does not add additional fragments. This is due to the reason that after sending the initial fragment, the daemon on name servers removes all RRs except RRSIG, DNSKEY, and OPT, thereby accommodating double-signatures created with SPHINCS+ in the same number of fragments as with SPHINCS+ alone. Mean resolution times with various combinations of double-signatures are shown in Tables 2, 3, 4. The results show that the impact of double signatures is insignificant on the resolution times of queries when compared with post-quantum signatures alone. In any double-signature combination that we have tested, resolution time only increases by less than 9% as compared with the post-quantum signatures alone. For instance, the resolution time for a combination of FALCON + ECDSA is 205.9ms as compared with 190.1ms for FALCON alone (i.e., an increase of approximately 8.3%). The lowest resolution time for double signatures was observed for any combination of FALCON with either RSA or ECDSA-based pre-quantum digital signatures. This is because FALCON has the shortest signature size compared with other post-quantum digital signatures. Similarly, RSA combined with any of the three post-quantum candidates results in a lower resolution time as compared with ECDSA (e.g., DILITHIUM + RSA results in a lower resolution time in comparison with DILITHIUM + ECDSA, and likewise for other two post-quantum digital signatures). This is because RSA has faster signature verification than ECDSA (Brady, 2020).
@@
The aforementioned observations suggest that double signatures can reliably be incorporated into DNNSEC infrastructure and have a minimal impact on performance as compared with post-quantum digital signatures alone. Therefore, their incorporation is ideally suited for DNSSEC during the interim period from now until CRQCs are fully realized. Particularly, the double signatures created with either FALCON or DILITHIUM with RSA or ECDSA have a resolution time that is 14-22% less than that of SPHINCS+ alone. Since both FALCON and DILITHIUM are based upon lattices, SPHINCS+ was only added as a post-quantum candidate by NIST to avoid the over-reliance on the lattices (see (NIST, 2022)). However, as observed in our experiments, SPHINCS+ has a significantly larger signature size, which results in much higher resolution time in DNSSEC. Therefore, the double-signature created with either FALCON or DILITHIUM is the optimal solution for the specified interim period while avoiding the envisaged problems of SPHINCS+ (see results described above and depicted in Tables 2 - 4). These double-signatures will essentially offer additional security to DNSSEC during the interim period, if any attacks on lattice-based signatures are revealed in the near future that could be perpetrated through today’s classical computer.
@@ 7. Related Work
Our literature review suggests that, historically, there have been a few proposals for dealing with large message size issues. For example, authors in (M. Sivaraman, 2015; Song and Wang., 2019) proposed potential ways for conducting fragmentation at the application layer. However, both of them can potentially lead to DoS amplification attacks since they send many packets in response to a single query. Additionally, there are a few recent research works that focus on transitioning toward quantum-safe DNSSEC and related issues. However, unlike this work, none of these prior research works consider double-signatures for DNSSEC. For example, authors in (Müller et al., 2020; Beernink, 2022; Goertzen and Stebila, 2022) investigated NIST Round 3 and 4 PQC candidates, respectively, for their utilization in DNSSEC. Müller et al. (Müller et al., 2020) specifically analyzed NIST Round 3 candidates and established the various requirements that different candidates either meet or do not for their incorporation in DNSSEC. Similarly, Beernink et al. (Beernink, 2022) presented a way of delivering larger keys out-of-band from DNSSEC. Goertzen et al. (Goertzen and Stebila, 2022) focused on the fragmentation issue in DNSSEC for the post-quantum era. However, as discussed before, it adopts a custom-designed non-standard pseudo-RR for performing the fragmentation at the application layer. The usage of non-standard RR makes its practicability questionable. Similarly, an improved version of the fragmentation approach for quantum-safe DNSSEC is also presented in (Rawat and Jhanwar, 2024), which leverages only standard RRs and thus seems more practical. However, unlike these existing works, we investigate the possibility of using double signatures for DNSSEC, and how the fragmentation issue can be dealt with in the simultaneous presence of pre- and post-quantum signatures in DNSSEC. Through our implemented testbed and empirical analysis, we show the efficacy of double signatures in DNSSEC for the interim period until CRQCs are fully realized.
@@ 8. Conclusion & Future Work
In this work, we investigated the possibility of using double signatures in DNSSEC. For this purpose, we first investigated the maximum message size issue resulting from the incorporation of double signatures in depth and then devised an approach that can successfully incorporate double signatures on name servers and deliver them to the resolver over UDP by conducting the fragmentation and re-assembly of response messages at the application layer. We then built a functional Docker-based testbed by leveraging OQS-BIND9 and a daemon that handles double signatures, fragmentation of relevant response messages, and the reassembly of the received fragments. Our investigation suggests double signatures can be generated and added to signed zone files on the authoritative name servers, and our resolver can facilitate the verification of two signatures for the same resource record. With these required changes made to the BIND9 resolver and clubbing our daemon for fragmentation/re-assembly of double-signatures with it, we comprehensively analyzed the resolution time for various combinations of pre- and post-quantum signatures. Our analysis revealed that double signatures do not have any significant overhead as compared with post-quantum digital signatures alone. Particularly, our investigation suggests that lattice-based signatures - i.e., FALCON and DILITHIUM combined with any of the pre-quantum digital signatures result in a shorter resolution time than SPHINCS+ alone, which was selected (despite having a very large signature size) by NIST to avoid over-reliance on the lattice-based signatures for the post-quantum era. Our analysis suggests that double signatures generated with FALCON and DILITHUM are well-suited for DNSSEC and offer performance benefits compared with SPHINCS+ alone. The double signatures approach will also offer additional security in case any attacks on lattice-based PQ candidates are revealed in the near future. While the analysis conducted in this work establishes the overall efficacy of double signatures for DNSSEC, we believe that resolution time can further be reduced by incorporating our fragmentation strategy directly into BIND9 rather than in the form of a daemon. We plan to make these advancements and relevant analyses in the future.